# Squish Jamming

### v1

@@ Abstract
A wide range of disordered materials, from biological to geological assemblies, feature discrete elements undergoing large shape changes. How significant geometrical variations at the microscopic scale affect the response of the assembly, in particular rigidity transitions, is an ongoing challenge in soft matter physics. However, the lack of a model granular-like experimental system featuring large and versatile particle deformability impedes advances. Here, we explore the oscillatory shear response of a sponge-like granular assembly composed of highly compressible elastic rings. We highlight a progressive rigidity transition, switching from a fluid-like to a solid-like response by increasing density or decreasing shear amplitude. The rearranging fluid state consists of crystal clusters separated by melted regions; in contrast, the solid state remains amorphous and absorbs all imposed shear elastically. We rationalise this transition by uncovering an effective, attractive shear force between rings that emerges from a friction-geometry interplay. If friction is sufficiently high compared to shear, the extent of the contacts between rings, captured analytically by elementary geometry, controls the rigidity transition.
@@
Particulate assemblies, such as granular media, can switch their mechanical behaviour between a solid and a fluid-like response depending on their density, applied stress or interactions. This solid-fluid transition is of tremendous importance across length scales and disciplines, with the elemental particles varying in size, shape, and stiffness. Such jamming or yielding transitions are, for instance, underlying the onset of geological flows [1], the transport and clogging in industrial food processing [2], and are at the heart of numerous biological processes [3].
@@
For rigid particles without internal degrees of freedom, the rigidity transition is mainly controlled by the average number of constraints imposed by the number of neighbours [4], including for non-spherical particles [5]. Even when particles are softer and can moderately deform their shape upon contact, such as emulsions or foams, the number of contacts remains an essential factor [6, 7, 8]. However, particle deformations may reach up to the order of the particle size for a broader class of materials, from biological tissues to porous soils. By adapting their shape to external constraints, such squishy particles [9] acquire additional degrees of freedom, potentially affecting the fluid-solid transition ingredients.
@@
Recent works highlight the role of shape and deformability in the jamming transition of highly deformable particles, both in thermal [10, 11] and athermal [12, 13, 14] assemblies. Dedicated tools have also been developed to quantify the large shape changes that cells undergo while embedded in a deforming tissue [15]. The shape of the cells in epithelial tissues is one of the essential parameters governing a rigidity transition [16, 17, 18]. In geological flows, the shape of constitutive particles undergoes significant changes primarily via breaking [19], which in turn have a critical effect on the rheology [20] and jamming transition [21, 1]. With very soft grains that do not break, compaction far above jamming highlights that not only the coordination number increases with compaction, but also the size of the contacts [22, 23].
@@
Despite these recent advances, a clear picture of how extreme deformability affects the fluid-solid transition is still lacking. It is an ongoing challenge in soft matter physics and related fields [24]. Although advanced numerical methods, such as the vertex model [18], the material point method [25] and phase field approach [26, 27], have been used, numerical investigations remain incredibly challenging owing to the non-linear couplings between geometry and interactions at the local scale, and a large number of elements. As such, a simple experimental model featuring large deformation at the grain scale would be essential to explore this question. One may leverage the recent advances in granular metamaterials [28] and metafluids [29], whose properties are governed by geometrical non-linearities at the particle scale, to control grains compressibility.
@@
Here, we introduce assemblies of compressible elastic rings as an experimental model to explore the role of large grain deformations in the rheology of disordered media. Via oscillatory shear experiments, we highlight a solid-fluid transition controlled by the shear amplitude and the density. While the transition is characterised by the absence or presence of structural rearrangements, compression-induced large geometrical deformations help the assembly absorb shear without rearranging in the solid phase. Finally, we argue that this transition is mainly driven by the size of the frictional contacts between the compressed rings via effective adhesive interactions under shear, which stem from the interplay between geometrical deformations and friction.
I Fluid-solid transition in a sheared assembly of compressible rings
@@
Figure 1: Fluid-solid transition under oscillatory shear of an assembly of compressible rings. a, Schematic of the experimental setup. A ring assembly is confined within a 2D cell with a fixed area. Oscillatory shear $\gamma=(\gamma_{0}/2)\sin\left(\dot{2\gamma_{0}}t/\gamma_{0}\right)$ is applied to the assembly via lateral motion of the top arm while the bottom one is fixed. Both top and bottom arms are decorated with half-solid disks of the same dimension as the rings to transmit shear. A transparent acrylic plate placed on top keeps the rings in the plane. b, Experimental and simulated (FEM) indentation response $f_{y}(\delta_{y})$, and c, associated stiffness $\frac{\mathrm{d}f_{y}}{\mathrm{d}\delta_{y}}$ of a single ring compressed between two flat plates. The experimental curve is the averaged response over five different rings from the assembly; upon loading and unloading, the shaded area shows the minimal and maximal values reached by the rings during the compression cycle. The corresponding FEM response is made via a 3D simulation of the neo-Hookean material model with Young’s modulus $E=1\,\mathrm{MPa}$ and Poisson’s ratio $\nu=0.45$, reproducing the ring and indentation geometry. The experiments and FEM, which agree quantitatively, feature a slight softening upon indentation, leading to solid foam-like deformation heterogeneities upon compression of the ring assembly. d, Top view of the ring assembly ($\Phi=1.04$ and $\gamma_{0}=0.53$) and representative of the pictures used for image analysis. Here, each tracked ring is coloured over its detected area by the value of its irreversible displacement $\log(\Delta u/r)$. The corresponding displacement is shown by a red arrow (up-scaled by a factor $4$). e, For each ring, we measure via image analysis its position (geometric centre), shape, neighbours, and extent of each contact (length of the blue segments). We define two neighbourhood types: Voronoi neighbours share a common edge in the Voronoi tessellation, and touching neighbours are in contact with a finite contact extent $\ell_{c}$. The thin orange silicone layer is applied in the inner part of the ring to help with detection and further imaging processes. See also Supplementary Video 2. f, Shear response $F_{X}(\gamma)$ of the assembly for $\Phi=1.00$ and for shear amplitude $\gamma_{0}$ ranging from $0.07$ to $0.7$. Six cycles in the steady state are displayed for each $\gamma_{0}$. g, Intensity of the averaged irreversible displacement $\log(\left\langle{\Delta u}\right\rangle/r)$ for varying density $\Phi$ and shear amplitude $\gamma_{0}$. The average is calculated over all tracked rings and in the steady state. h, Map of the ratio between the loss and storage moduli of the assembly, computed from the shear response $F_{x}(\gamma)$ and averaged over the cycles in the steady state. The corresponding $F_{x}(\gamma)$ for $\Phi=1.00$ are shown in f, each coloured by its corresponding value of $\log(\langle G^{\prime\prime}\rangle/\langle G^{\prime}\rangle)$. In g,h, and subsequent maps, the white dotted line is a geometric prediction introduced thereafter (Eq. (1)), but now serves as a reference to help compare the maps and identify solid and fluid regions.
@@
We consider a two-dimensional assembly of frictional, elastic, and ring-shaped grains (see Fig. 1a and Supplementary Video 1). The assembly is compressed into an imposed area and sheared by oscillatory displacement of the top boundary as shown in the setup schematic in Fig. 1a. The rings are all identical with a thickness $t=1.5\,\mathrm{mm}$, radius $r=3.3\,\mathrm{mm}$ and height $h=10\,\mathrm{mm}$. Rings are fabricated via moulding of polymerising liquid silicone (see Methods and Supplementary Fig. LABEL:S-fig_sup_ring), then coated with talc powder to avoid adhesion and ensure dry frictional interactions. The compression response of the rings to a loading-unloading cycle (Fig. 1b) attests to a perfect elasticity with no remaining permanent deformations and high reproducibility, and is well recovered by the finite element method (FEM) (see Methods for details). In addition, the onset of the compression response features a small stiffness decrease (Fig. 1c). Deformed rings being slightly softer, a compressed assembly of such rings develops heterogeneous deformations as shown in the top view displayed in Fig. 1d (see also Supplementary Video 1). This heterogeneous deformation upon compression is reminiscent of the mechanics of cellular materials [30, 31, 32], including solid foams [33], with the critical difference that here the unit cells do not adhere each other and can therefore rearrange.
@@
The assembly’s response to shear is explored by systematically varying the shear amplitude $\gamma_{0}$ and the density of rings $\Phi$. The oscillatory shear $\gamma$ is imposed at a fixed maximum shear rate of $\dot{\gamma_{0}}=1.33\times 10^{-2}\,\mathrm{s^{-1}}$, low enough to ensure a quasi-static regime where inertial or vibrational effects are negligible. The density is adjusted by varying the number of rings in the shear cell and quantified by $\Phi=N\pi r_{h}^{2}/A$ with $r_{h}=r+t/2$ the external radius of the rings, $N$ the number of rings and $A$ the area of the shear cell. This is the density that would correspond to the undeformed area of the rings, and we focus on the regime above the hexagonal packing density $\Phi_{h}=\frac{\pi}{2\sqrt{3}}\approx 0.91$. Initialisation is made by dispersing the rings in the cell area and then manually lowering and fixing the top arm at the desired height. The ring assembly is, therefore, initially in a mechanically quenched state. An acrylic plate on top of the assembly holds the rings in 2D, preventing out-of-plane buckling.
@@
The mechanical response to shear is probed locally via image analysis and globally via a force sensor. Pictures from an overhanging camera are taken regularly during the cycles, such that the position, shape and neighbours of each ring are identified and tracked all along the shear cycles (Fig. 1e; see also Methods and Supplementary Video 2). A force sensor, placed between the driving screw and the top arm, records the force $F_{X}$ necessary to shear the assembly. The shear force $F_{X}$ for $\Phi=1.0$ and all investigated shear amplitudes is shown in Fig. 1f. Unless stated otherwise, the presented measures represent the system in the steady state (see Supplementary Table LABEL:S-suptab1 for the cycles used in the analysis). The high reproducibility of the shear response in Fig. 1f testifies that the assembly indeed reached a steady state.
@@
We then assess whether the ring assembly responds more in a solid-like (jammed) or fluid-like (yielded) fashion. To assess this response, a classic measure for particulate systems under oscillatory loading is the irreversible displacement of the particles between two consecutive cycles [34, 35, 36, 37, 38, 39]. We define $\Delta u(i,c)$ as the norm of the irreversible displacement of ring $i$ occurring between the start and end of cycle $c$. In Fig. 1d, rings are coloured by their value of $\log\left(\Delta u/r\right)$ and the associated irreversible displacement is drawn with a red arrow (magnified by a factor 4). Even at such a high density, irreversible displacements occur but are highly heterogeneous, with values spanning orders of magnitude within the assembly. To explore how the irreversibility depends on the density and shear amplitude, for each parameter pair $(\Phi,\gamma_{0})$ considered, we compute $\langle\Delta u\rangle$, the averaged irreversible displacement over all tracked rings in the steady state. The resulting map is shown in Fig. 1g. Two zones can be identified: a fluid-like regime with significant irreversibility, which smoothly transitions to a solid-like zone with barely any irreversible displacement, by increasing density or decreasing shear (see also non-affine trajectories in Supplementary Video 3).
@@
The transition between the fluid-like and solid-like features is observed not only in the rings’ local movement but also via the global mechanical response. From the shear response $F_{X}(\gamma)$ of the whole assembly, we compute for each cycle the first harmonic of the storage and loss moduli, $G^{\prime}$ and $G^{\prime\prime}$, respectively [40] (see separated map of $G^{\prime\prime}$ and $G^{\prime}$ in Extended Data Fig. 1). Identifying when the ratio $G^{\prime\prime}/G^{\prime}$ is greater than $1$ is a standard method to evaluate if the assembly yields or remains essentially solid [41]. The map of $\log(\left\langle{G^{\prime\prime}}\right\rangle/\left\langle{G^{\prime}}\right\rangle)$ shown in Fig. 1h, with $\langle{G^{\prime\prime}}\rangle,\langle{G^{\prime}}\rangle$ the moduli averaged over the cycles in the steady state, recovers qualitatively the solid- and fluid-like regions identified above from the irreversible displacements (Fig. 1g). Indeed, the region with high irreversible displacements coincides with a loss modulus significantly larger than the storage one and inversely for the dominantly reversible region.
@@
Exploring the mechanical response to shear at very high density and shear amplitude, we identify a progressive jamming transition of the ring assembly by either reducing the shear amplitude or increasing the density. To understand the driving factors of this transition, we first quantify the structural properties of the assembly.
II A partially crystallised fluid and an amorphous solid
Figure 2: Crystallising fluid and amorphous solid. a, Left, probability density profiles of the hexagonal bond orientational order $\Psi_{6}$ for an assembly at $(\gamma_{0},\Phi)=(0.4,1.00)$ in the yielded region. The profiles are shown for the assembly before shear and after $20$, $40$ and $60$ shear cycles. Right, corresponding snapshots of the assembly after cycling, where each ring is coloured by its $\Psi_{6}$ value. The assembly shows a progressive growth of crystalline regions separated by melted ones. b, Same quantities as in a but for an assembly with higher density and same shear amplitude, $(\gamma_{0},\Phi)=(0.4,1.04)$, now located in the jammed region. Here, no significant structural changes are observed, and despite high shear, the assembly remains amorphous. See also Supplementary Video 4. c, Map of the proportion of rings belonging to crystal clusters (with $\Psi_{6}>0.75$) in the steady state. The corresponding snapshots of the assemblies in their steady states, with the rings coloured by their $\Psi_{6}$ values, are shown on the right for $\gamma_{0}=0.25$ and all investigated densities. See also Supplementary Video 5 for the time evolution of $\Psi_{6}$.
@@ II A partially crystallised fluid and an amorphous solid
The presence or absence of irreversible displacements upon cyclic shear in the steady state highlights that ring rearrangements and associated structural deformations play a crucial role in the mechanical response of the assembly [42, 43]. We then measure the structural differences between the solid and fluid regions and how they evolve from their initial quenched state. With all rings having the same size, the configuration with the highest density and minimal shape deformation (and hence minimum elastic energy) is the hexagonal packing. We thus use the hexagonal bond orientational order $\Psi_{6}$ as the structural indicator of the assembly [44]. For each ring $j$, the hexagonal order is computed by $\Psi_{6}(j)=(1/n_{j})|{\sum_{k=1}^{n_{j}}\mathrm{e}^{6\mathrm{i}\alpha_{jk}}}|$ with $n_{j}$ the number of Voronoi neighbours and $\alpha_{jk}$ the angle between the line passing through the centre of the two rings and the $x$ axis (the direction of the shear). The value of $\Psi_{6}$ is $1$ for a ring in a perfect hexagonal lattice and continuously decreases to 00 as the structure deviates from the hexagonal one.
@@
As the assembly accumulates shear cycles, we compare in Fig. 2 (see also Supplementary Video 4) the structural evolution of configurations for two cases located near the border between the jammed and yielded regions, with each case positioned on the opposite side of the border. After mechanical quenching, both configurations start in an amorphous state with a flat probability distribution $P(\Psi_{6})$, but their evolution with cycling strongly differ. In the yielded region, for $(\gamma_{0},\Phi)=(0.4,1.00)$ shown in Fig. 2a, cycling induces drastic structural changes with the progressive growth of distinct crystal clusters. The probability density increases close to $\Psi_{6}=1$ while it decreases for low values. In contrast, for larger densities within the jammed region (with $(\gamma_{0},\Phi)=(0.4,1.06)$ shown in Fig. 2b), no significant structural changes occur upon cycling, and the assembly remains in an amorphous phase. We then measure the crystallised portion of the assembly, or proportion of rings having $\Psi_{6}>0.75$, in the steady state over the parameter space (Fig. 2c). All the assemblies in the solid region feature almost no structural changes and remain amorphous, while the proportion of the crystal clusters decreases with increasing density in the fluid region (Fig. 2c; see also Supplementary Video 5). Rearranging rings within a given surface area tend to crystallise to reduce their overall elastic energy, as the hexagonal packing is the most compact structure without shape deformations. However, the growth of crystal clusters is limited by the presence of rings deformed by compression, which necessarily exist for high densities and prevent crystallisation.
@@
The structure of the ring assembly in its steady state differs drastically between the yielded and jammed regions. The fluid-like region is characterised by hexagonal crystal clusters separated by a melted, disordered phase. This partial crystallisation is typical of granular media under sufficient shear where reorganisations lead to structural changes [45, 46, 47, 48]. Even if essentially crystallised, the yielded phase still displays a fluid-like mechanical response with irreversible displacements (Fig. 1g,h). In contrast, no crystallisation occurs in the jammed region, where the assembly remains amorphous almost without rearrangements. Even under high shear, the rings stick together to keep the same structure. In the following, we delve into the dynamics during shear cycles to investigate the interplay between structural and geometrical deformations of the rings.
III Large geometrical deformations absorb shear and hinder rearrangements
Figure 3: Competition between geometrical and structural deformations. a, The geometrical deformation of the rings is quantified by the shape index $q$. The evolution of $q$ within a cycle is shown for one ring in a $(\gamma_{0},\Phi)=(0.4,1.00)$ assembly. Each marker goes with an associated ring snapshot. We extract $\tilde{q}$ and $\Delta q$, the average and amplitude variation from the shape trajectory. b,c, Average shape of the assembly $\tilde{q}$ and shape variation $\Delta q$ as a function of shear amplitude for each density. The values are averaged over all tracked rings in the steady state. d,e, Map of the geometrical $S_{c}$ and topological $T_{c}$ strains of the assembly. The total strains here are the cumulative sums of the norm of the frame-to-frame strains estimated over the whole assembly and in the steady state.

### v2

@@ Abstract
A wide range of disordered materials, from biological to geological assemblies, feature discrete elements undergoing large shape changes. How significant geometrical variations at the microscopic scale affect the response of the assembly, in particular rigidity transitions, is an ongoing challenge in soft matter physics. However, the lack of a model granular-like experimental system featuring large and versatile particle deformability impedes advances. Here, we explore the oscillatory shear response of a sponge-like granular assembly composed of highly compressible elastic rings. We highlight a progressive rigidity transition, switching from a yielded phase to a solid one by increasing density or decreasing shear amplitude. The rearranging yielded state consists of crystal clusters separated by melted regions; in contrast, the solid state remains amorphous and absorbs all imposed shear elastically. We rationalize this transition by uncovering an effective, attractive shear force between rings that emerges from a friction-geometry interplay. If friction is sufficiently high, the extent of the contacts between rings, captured analytically by elementary geometry, controls the rigidity transition.
keywords:
rigidity transition |||| disordered media |||| geometry |||| friction
[…]
This manuscript was compiled on December 13, 2024 www.pnas.org/cgi/doi/10.1073/pnas.XXXXXXXXXX
@@
Particulate assemblies, such as granular media, can switch their mechanical behavior between a solid and a fluidized response depending on their density, applied stress or interactions. This rigidity transition is of tremendous importance across length scales and disciplines, with the elemental particles varying in size, shape, and stiffness. Such jamming or yielding transitions are, for instance, underlying the onset of geological flows (1), the transport and clogging in industrial food processing (2), and are at the heart of numerous biological processes (3).
@@
For athermal rigid particles without internal degrees of freedom, the rigidity transition, called jamming, is mainly controlled by the average number of constraints imposed by the number of neighbors (4), including for non-spherical particles (5). Even when particles are softer and can moderately deform their shape upon contact, such as emulsions or foams, the number of contacts remains an essential factor (6, 7, 8). However, particle deformations may reach up to the order of the particle size for a broader class of materials, from biological tissues to porous soils. By adapting their shape to external constraints, such "squishy" particles (9) acquire additional degrees of freedom, potentially affecting the ingredients underlying the jamming transition.
@@
Recent works highlight the role of shape and deformability in the jamming transition of highly deformable particles, both in thermal (10, 11) and athermal (12, 13, 14, 15) assemblies. Dedicated tools have also been developed to quantify the large shape changes that cells undergo while embedded in a deforming tissue (16). The shape of the cells in epithelial tissues is one of the essential parameters governing a rigidity transition (17, 18, 19, 20). In geological flows, the shape of constitutive particles undergoes significant changes primarily via breaking (21), which in turn have a critical effect on the rheology (22) and jamming transition (23, 1). With very soft grains that do not break, compaction far above jamming highlights that not only the coordination number increases with compaction, but also the size of the contacts (24, 25).
@@
Despite these recent advances, a clear picture of how extreme deformability affects the jamming transition is still lacking. It is an ongoing challenge in soft matter physics and related fields (26). Although advanced numerical methods, such as the vertex model (19), the material point method (27) and the phase field approach (28, 29), have been used, numerical investigations remain incredibly challenging owing to the non-linear couplings between geometry and interactions at the local scale, and a large number of elements. As such, a simple experimental model featuring large deformation at the grain scale would be essential to explore this question. One may leverage the recent advances in granular metamaterials (30) and metafluids (31), whose properties are governed by geometrical non-linearities at the particle scale, to control grains compressibility.
@@
Here, we introduce assemblies of compressible elastic rings as an experimental model to explore the role of large grain deformations in the rheology of disordered media. Via oscillatory shear experiments, we highlight a rigidity transition controlled by the shear amplitude and the density. While the transition is characterized by the absence or presence of structural rearrangements, compression-induced large geometrical deformations help the assembly absorb shear without rearranging in the solid phase. Finally, we argue that this transition is mainly driven by the size of the frictional contacts between the compressed rings via effective adhesive interactions under shear, which stem from the interplay between geometrical deformations and friction.
Rigidity transition in a sheared assembly of compressible rings
@@
Figure 1: Oscillatory shear of an assembly of compressible rings. (A) Schematic of the experimental setup. A ring assembly is confined within a 2D cell with a fixed area. Oscillatory shear $\gamma=(\gamma_{0}/2)\sin\left(\dot{2\gamma_{0}}t/\gamma_{0}\right)$ is applied to the assembly via lateral motion of the top arm while the bottom one is fixed. Both top and bottom arms are decorated with half-solid disks of the same dimension as the rings to transmit shear. A transparent acrylic plate placed on top keeps the rings in the plane. The inset shows an image of a single ring. The thin orange silicone layer is applied in the inner part of the ring to help with detection and further imaging processes. (B,C) Experimental and simulated (FEM) indentation response $f_{y}(\delta_{y})$ (B) and associated stiffness $\frac{\mathrm{d}f_{y}}{\mathrm{d}\delta_{y}}$ (C) of a single ring compressed between two flat plates. The experimental curve is the averaged response over five different rings from the assembly; upon loading and unloading, the shaded area shows the minimal and maximal values reached by the rings during the compression cycle. The corresponding FEM response is made via a 3D simulation of the neo-Hookean material model with Young’s modulus $E=1\,\mathrm{MPa}$ and Poisson’s ratio $\nu=0.45$, reproducing the ring and indentation geometry. The experiments and FEM, which agree quantitatively, feature a slight softening upon indentation, leading to solid foam-like deformation heterogeneities upon compression of the ring assembly. (D) For each ring, we measure via image analysis its position (geometric center), shape, neighbors, and extent of each contact (length of the blue segments). We define two neighborhood types: Voronoi neighbors share a common edge in the Voronoi tessellation, and touching neighbors are in contact with a finite contact extent $\ell_{c}$. The bond vectors between touching neighbors are represented by dashed arrows. See also Movie S2.
@@
We consider a two-dimensional assembly of frictional, elastic, and ring-shaped grains (see Fig. 1A and Movie S1). To explore the rheological response of the ring assembly, we use a typical oscillatory shear method (32, 33, 34, 35, 36) where the assembly is compressed into an imposed area and sheared by oscillatory displacement of the top boundary as shown in the setup schematic in Fig. 1A. The rings are all identical with a thickness $t=1.5\,\mathrm{mm}$, radius $r=3.3\,\mathrm{mm}$ and height $h=10\,\mathrm{mm}$. Rings are fabricated via molding of polymerizing liquid silicone (see Methods and SI Appendix, Fig. LABEL:S-fig_sup_ring), then coated with talc powder to avoid adhesion and ensure dry frictional interactions. The compression response of the rings to a loading-unloading cycle (Fig. 1B) attests to a perfect elasticity with no remaining permanent deformations and high reproducibility, and is well recovered by the finite element method (FEM) (see Methods for details). In addition, the onset of the compression response features a small stiffness decrease (Fig. 1C). Deformed rings being slightly softer, a compressed assembly of such rings develops heterogeneous deformations as shown in the top view displayed in Fig. 1D (see also Movie S1). This heterogeneous deformation upon compression is reminiscent of the mechanics of cellular materials (37, 38, 39), including solid foams (40), with the critical difference that here the unit cells do not adhere each other and can therefore rearrange.
@@
The assembly’s response to shear is explored by systematically varying the shear amplitude $\gamma_{0}$ and the density of rings $\Phi$. The oscillatory shear $\gamma$ is imposed at a fixed maximum shear rate of $\dot{\gamma_{0}}=1.33\times 10^{-2}\,\mathrm{s^{-1}}$, low enough to ensure a quasi-static regime where inertial or vibrational effects are negligible. The density is adjusted by varying the number of rings in the shear cell and quantified by $\Phi=N\pi r_{h}^{2}/A$ with $r_{h}=r+t/2$ the external radius of the rings, $N$ the number of rings and $A$ the area of the shear cell. This is the density that would correspond to the undeformed area of the rings, and we focus on the regime above the hexagonal packing density $\Phi_{h}=\frac{\pi}{2\sqrt{3}}\approx 0.91$. Initialization is made by dispersing the rings in the cell area and then manually lowering and fixing the top arm at the desired height. The ring assembly is, therefore, initially in a mechanically quenched state. An acrylic plate on top of the assembly holds the rings in 2D, preventing out-of-plane buckling (41).
@@
The mechanical response to shear is probed locally via image analysis and globally via a force sensor. Pictures from an overhanging camera are taken regularly during the cycles, such that the position, shape and neighbors of each ring are identified and tracked all along the shear cycles (Fig. 1D; see also Methods and Movie S2). Image analysis permits to measure quantities such as the irreversible displacement between two consecutive cycles (Fig. 2A). A force sensor, placed between the driving screw and the top arm, records the force $F_{X}$ necessary to shear the assembly. The shear force $F_{X}$ for $\Phi=1.0$ and all investigated shear amplitudes is shown in Fig. 2B. Unless stated otherwise, the presented measures represent the system in the steady state (see SI Appendix, Table LABEL:S-suptab1 for the cycles used in the analysis). The high reproducibility of the shear response in Fig. 2B testifies that the assembly indeed reached a steady state and that results are robust against which cycles are considered in the steady state (SI Appendix, Fig. LABEL:S-fig_sup_same_cycle). We also verified that in the steady state, the experiment is reproducible and not sensitive to the initial configuration (SI Appendix, Fig. LABEL:S-fig_sup_repro).
Figure 2: Rigidity transition in the density-shear amplitude phase space. (A) Top view of a portion of the ring assembly ($\Phi=1.04$ and $\gamma_{0}=0.53$). Here, each tracked ring is colored over its detected area by the value of its irreversible displacement $\log(\Delta u/r)$. The corresponding displacement is shown by a red arrow (up-scaled by a factor $4$). (B) Shear response $F_{X}(\gamma)$ of the assembly for $\Phi=1.00$ and for shear amplitude $\gamma_{0}$ ranging from $0.07$ to $0.7$. Six cycles in the steady state are displayed for each $\gamma_{0}$. (C) Intensity of the averaged irreversible displacement $\log(\left\langle{\Delta u}\right\rangle/r)$ for varying density $\Phi$ and shear amplitude $\gamma_{0}$. The average is calculated over all tracked rings and in the steady state. (D) Map of the ratio between the loss and storage moduli of the assembly, computed from the shear response $F_{x}(\gamma)$ and averaged over the cycles in the steady state. The corresponding $F_{x}(\gamma)$ for $\Phi=1.00$ are shown in (B), each colored by its corresponding value of $\log(\langle G^{\prime\prime}\rangle/\langle G^{\prime}\rangle)$. In ($C$ and $D$) and subsequent maps, the white dotted line is a geometric prediction introduced later (Eq. [1]), which serves here as a reference to help compare the maps and identify solid and yielded regions.
@@
We then assess whether the ring assembly responds more in a solid-like or yielded fashion. To assess this response, a classic measure for particulate systems under oscillatory loading is the irreversible displacement of the particles between two consecutive cycles (42, 43, 44, 34, 35, 36, 45). We define $\Delta u(i,c)$ as the norm of the irreversible displacement of ring $i$ occurring between the start and end of cycle $c$. In Fig. 2A, rings are colored by their value of $\log\left(\Delta u/r\right)$ and the associated irreversible displacement is drawn with a red arrow (magnified by a factor 4). Even at such a high density, irreversible displacements occur but are highly heterogeneous, with values spanning orders of magnitude within the assembly. To explore how the irreversibility depends on the density and shear amplitude, for each parameter pair $(\Phi,\gamma_{0})$ considered, we compute $\langle\Delta u\rangle$, the averaged irreversible displacement over all tracked rings in the steady state. The resulting map is shown in Fig. 2C. Two zones can be identified: a yielded regime with significant irreversibility, which smoothly transitions to a solid-like zone with barely any irreversible displacement, by increasing density or decreasing shear (see also non-affine trajectories in Movie S3).
@@
The transition between the yielded and solid-like features is observed not only in the rings’ local movement but also via the global mechanical response. From the shear response $F_{X}(\gamma)$ of the whole assembly, we compute for each cycle the first harmonic of the storage and loss moduli, $G^{\prime}$ and $G^{\prime\prime}$, respectively (46) (see separated map of $G^{\prime\prime}$ and $G^{\prime}$ in SI Appendix, Fig. LABEL:S-fig_sup_moduli). Identifying when the ratio $G^{\prime\prime}/G^{\prime}$ is greater than $1$ is a standard method to evaluate if the assembly yields or remains essentially solid (47). The map of $\log(\left\langle{G^{\prime\prime}}\right\rangle/\left\langle{G^{\prime}}\right\rangle)$ shown in Fig. 2D, with $\langle{G^{\prime\prime}}\rangle,\langle{G^{\prime}}\rangle$ the moduli averaged over the cycles in the steady state, recovers qualitatively the solid and yielded regions identified above from the irreversible displacements (Fig. 2C). Indeed, the region with high irreversible displacements coincides with a loss modulus significantly larger than the storage one and inversely for the dominantly reversible region.
@@
Exploring the mechanical response to shear at very high density and shear amplitude, we identify a progressive rigidity transition of the ring assembly by either reducing the shear amplitude or increasing the density. To understand the driving factors of this transition, we first quantify the structural properties of the assembly.
A partially crystallized yielded assembly and an amorphous solid
@@
The presence or absence of irreversible displacements upon cyclic shear in the steady state highlights that ring rearrangements and associated structural deformations play a crucial role in the mechanical response of the assembly (48, 49). We then measure the structural differences between the solid and yielded regions and how they evolve from their initial quenched state. With all rings having the same size, the configuration with the highest density and minimal shape deformation (and hence minimum elastic energy) is the hexagonal packing. We thus use the hexagonal bond orientational order $\Psi_{6}$ as the structural indicator of the assembly (50). For each ring $j$, the hexagonal order is computed by $\Psi_{6}(j)=(1/n_{j})|{\sum_{k=1}^{n_{j}}\mathrm{e}^{6\mathrm{i}\alpha_{jk}}}|$ with $n_{j}$ the number of Voronoi neighbors and $\alpha_{jk}$ the angle between the line passing through the center of the two rings and the $x$ axis (the direction of the shear). The value of $\Psi_{6}$ is $1$ for a ring in a perfect hexagonal lattice and continuously decreases to 00 as the structure deviates from the hexagonal one.
@@
As the assembly accumulates shear cycles, we compare in Fig. 3 (see also Movie S4) the structural evolution of configurations for two cases located near the border between the solid and yielded regions, with each case positioned on the opposite side of the border. After mechanical quenching, both configurations start in an amorphous state with a flat probability distribution $P(\Psi_{6})$, but their evolution with cycling strongly differ. In the yielded region, for $(\gamma_{0},\Phi)=(0.4,1.00)$ shown in Fig. 3A, cycling induces drastic structural changes with the progressive growth of distinct crystal clusters. The probability density increases close to $\Psi_{6}=1$ while it decreases for low values. In contrast, for larger densities within the rigid region (with $(\gamma_{0},\Phi)=(0.4,1.06)$ shown in Fig. 3B), no significant structural changes occur upon cycling, and the assembly remains in an amorphous phase. We then measure the crystallized portion of the assembly, or proportion of rings having $\Psi_{6}>0.75$, in the steady state over the parameter space (Fig. 3C). All the assemblies in the solid region feature almost no structural changes and remain amorphous, while the proportion of the crystal clusters decreases with increasing density in the yielded region (Fig. 3C; see also Movie S5). Rearranging rings within a given surface area tend to crystallize to reduce their overall elastic energy, as the hexagonal packing is the most compact structure without shape deformations. However, the growth of crystal clusters is limited by the presence of rings deformed by compression, which necessarily exist for high densities and prevent crystallization.
Figure 3: Crystallizing yielded assembly and amorphous solid. (A) Left, probability density profiles of the hexagonal bond orientational order $\Psi_{6}$ for an assembly at $(\gamma_{0},\Phi)=(0.4,1.00)$ in the yielded region. The profiles are shown for the assembly before shear and after $20$, $40$ and $60$ shear cycles. Right, corresponding snapshots of the assembly after cycling, where each ring is colored by its $\Psi_{6}$ value. The assembly shows a progressive growth of crystalline regions separated by melted ones. (B) Same quantities as in (A) but for an assembly with higher density and same shear amplitude, $(\gamma_{0},\Phi)=(0.4,1.04)$, now located in the solid region. Here, no significant structural changes are observed, and despite high shear, the assembly remains amorphous. See also Movie S4. (C) Map of the proportion of rings belonging to crystal clusters (with $\Psi_{6}>0.75$) in the steady state. The corresponding snapshots of the assemblies in their steady states, with the rings colored by their $\Psi_{6}$ values, are shown on the right for $\gamma_{0}=0.25$ and varying densities. See also Movie S5 for the time evolution of $\Psi_{6}$.
@@
The structure of the ring assembly in its steady state differs drastically between the yielded and solid regions. The yielded region is characterized by hexagonal crystal clusters separated by a melted, disordered phase. This partial crystallization is typical of granular media under sufficient shear where reorganizations lead to structural changes (51, 52, 53, 54). Even if essentially crystallized, the yielded phase still displays significant plastic flow with irreversible displacements (Fig. 2C and $D$). In contrast, no crystallization occurs in the solid region, where the assembly remains amorphous almost without rearrangements. Even under high shear, the rings stick together to keep the same structure. In the following, we delve into the dynamics during shear cycles to investigate the interplay between structural and geometrical deformations of the rings.
Large geometrical deformations absorb shear and hinder rearrangements